OpenStudy (anonymous):

Lakewood Middle School is sponsoring a bus trip to the zoo. In order to pay for the bus at least 30 students must sign up for the trip. The bus can hold 68 students. Write an inequality to model this situation. a. 30x>68 c. 30x>68

OpenStudy (anonymous):

wouldnt it be c?

OpenStudy (anonymous):

c. 30<x<68 d. 30>x>68 supposed to be written as \[30 \le x \le68 \] & \[30 \ge x \ge 68\]

OpenStudy (anonymous):

it should be answer a

OpenStudy (anonymous):

it was , your right.. thanks

OpenStudy (anonymous):

answer c also correct since it is equal wth a. :-)